In a baseball team, 2 are Canadian, 3 are Americans, 1 is indian, and 3 are koreans. If a person is selected at random, what is the probability that the person is either canadian or korean?


1
Expert's answer
2021-04-15T06:33:58-0400
"P(canadian \ or \ korean)=\dfrac{2+3}{2+3+1+3}=\dfrac{5}{9}"